# Artificial Intelligence HCI Human-Computer Interaction

### v1

@@ 3.1. Study Materials
Our study consisted of a pre-questionnaire and a semi-structured interview including a sketching task. Our interview guide is available in Appendix A. Both the complete pre-questionnaire and the interview guide are also provided in our anonymized OSF repository.
@@
The interview guide was organized into three parts: (a) participants’ real-life QDA workflows (b) their pain points and gain points in the QDA process (c) their perspectives on the potential for AI support
@@
Lastly (c), we shifted to participants’ expectations regarding AI-assisted QDA. We inquired about the conditions under which they would be willing to integrate AI into their workflows, which steps they believed could be automated, and which aspects should remain human-driven. Finally, participants were given the opportunity to share any additional thoughts or insights on the topic.
@@ 3.2. Recruitment & Participants
We employed purposive and snowball sampling methods (Palinkas et al., 2015) to recruit participants who had conducted at least one study involving QDA through our professional networks. We approached researchers directly and invited them to refer interested colleagues, leveraging peer identification to assemble a diverse and well-qualified group. To capture a broad range of experiences and perspectives, we took care to recruit participants from diverse group affiliations, different study approaches, topic choices, and educational backgrounds. Our goal was to achieve data saturation regarding participants’ workflows, pain points, and gain points. This allowed us to identify common themes and insights across different stages of the QDA process, providing a solid foundation for our analysis.
@@
Table 1 provides a detailed overview of our participants. Our sample consisted of 15 HCI researchers with different backgrounds, coming from computer science, cybersecurity, engineering, and social sciences, now all firmly embedded in HCI research. Twelve participants were early career researchers (most of them working towards their PhD), and three senior researchers. Eight of them had worked on 1-3 QDA projects, four on 4-6, and three on more than 7, with years of experience in working with qualitative data varying from 1 to 9 years. We recruited participants until data saturation was reached. We achieved saturation around interview 12, and interviewed 3 more participants to confirm this.
Table 1. Detailed information on our interview participants’ research backgrounds, their QDA experience, and their primary QDA tools and methods.
@@
3.3. Study Procedure
Participants first completed the pre-questionnaire, which included an overview of the study procedure and a consent form. After that, we scheduled an appointment for the interview. The interviews were semi-structured, conducted face-to-face, and lasted approximately 45 minutes. Depending on the participant’s native language, six interviews were conducted in German and nine in English.. As the interviews were not recorded, two researchers were present: the primary interviewer, who led the session and took notes, and a second researcher, who created detailed protocols, sometimes verbatim. We decided not to record the interviews for the following reasons: First, to protect our participants’ privacy by adhering to data minimization principles (i. e., not collecting more data than necessary), especially since participants discussed recent research projects, some of which had not yet been published. Sensitive details about research content were intentionally excluded from notes. Second, to foster an informal and conversational atmosphere, as prior research has shown that recording can introduce stress (Leite et al., 1999; Noack et al., 2019) Moreover, we wanted participants to be open about their methodological choices, including possible shortcuts or subpar practices they may have taken, which they may be reluctant to share in case it jeopardizes their professional status or reputation. Thus, by not recording, we encouraged participants to share more about themselves. The primary interviewer, fluent in both German and English, conducted all interviews, ensuring consistency across sessions. The second researcher varied among four researchers and was either fluent in English or fluent in both German and English for German-language interviews. Immediately after each interview, the two researchers reflected on the session and synthesized their notes and protocols to capture accurate reflections of their attention rather than relying on memory. To ensure objectivity, they noted down points that stood out without performing any immediate analysis or interpretation. The researchers then compared, merged, and verified the completeness of their notes, ensuring no conflicts arose. Finally, when necessary, the researchers manually translated their synthesized interview protocols into English for subsequent data analysis.
We piloted our study with three participants to test the workflow and ensure the appropriateness of questions and their order. As no changes were made to the study procedure, the pilots were included in the final dataset, resulting in a sample size of 15 participants.
@@ 3.4. Ethical Considerations & Data Protection
We designed our study with strict ethical considerations according to the Menlo report (Kenneally and Dittrich, 2012). We obtained an ethical review and approval from our institution’s ethics committee for all study-related materials and a document outlining the planned study process. Before entering our study, all participants agreed to participate in a consent form attached to our pre-questionnaire. Before the interviews began, the interviewer walked the participants through the consent form again and gave them the opportunity to ask questions and asked for their consent again. To protect the privacy of the participants, they were given pseudonyms after completing the questionnaire and all data was pseudonymized. We stored and managed all data according to national data protection regulations and in accordance with our institution’s policies on confidentiality and data protection.
@@ 3.5. Data Analysis
Three authors prepared an initial set of deductive codes for tools and resources, data types, and participants’ overall willingness to use AI, based on the pre-questionnaire answer options and interview guide. They individually coded three protocols each for the coding process, pain points, fun points, and concerns and conditions for using AI within QDA to derive common themes. This process involved closely examining the protocols to identify and conceptualize subcategories. All three coders then created the final codebooks by discussing the identified themes and resolving conflicts in a collaborative session. On this basis, two of these authors coded all interview transcripts and discussed and resolved any disagreements to achieve complete agreement. To help us later illustrate our findings, the coders included example quotations from the protocols in our codebooks, some of which are verbatim and some paraphrased.
Additionally, we analyzed the drawings from the sketching task to derive commonalities and differences in the coding process and extracted participants’ suggestions for concrete AI applications as in vivo codes. Two coders jointly inspected both sketches and protocols, aiming to create a combined metaview over current QDA practices. First, each participant’s workflow was sketched out in a drawing. The replications were then sequentially mapped onto the meta-view. While new process elements were added to the design, already included elements were skipped. In the final step, we merged closely related processes to reduce the complexity of multiple single elements in the metaview. We differentiated elements taking into account the SeeMe-method developed by Herrmann (Herrmann, 2012). SeeMe is a semi-structured socio-technical modeling technique used for task analysis that involves three visualization elements: (1) roles assigned to individuals or teams, (2) activities that are of a more dynamic characteristic and represent change functions, and (3) entities, defined as objects from the physical world. For entities, we excluded tools and software support from the metaview as they were heavily dependent on the participant’s individual preference, but generally involved in each process element.
@@ 4. Results
Here, we present HCI researchers’ QDA workflows, including their pain points (Section 4.1) and report participants’ willingness to use AI support as well as their concerns and conditions (Section 4.2). In Appendix B, we provide the complete codebooks from our interviews, including descriptions, examples, and quantities for each code. Appendix C shows examples of participants’ workflow sketches.
@@ 4.1.1. Contexts of Studied QDA Workflows
In the pre-questionnaire, participants reported having analyzed open-ended survey responses ($n=12$), interview transcripts ($n=11$), written documents ($n=8$), and social media posts ($n=5$) as the four most frequent data types. When asked about the primary software used for qualitative coding, most participants ($n=9$) reported using more than one tool and specifically named, MAXQDA ($n=12$), Excel ($n=6$), and Google Docs ($n=4$). Participants also indicated experience with multiple methodological frameworks, and choosing among as well as adapting them based on the nature of the data and the specific objectives of each research project. Thematic Analysis ($n=11$), Grounded Theory ($n=6$), and Content Analysis ($n=4$) were mentioned most frequently.
@@ 4.1.2. Uncovering Participants’ Actual Coding Workflow
Figure 1. Detailed overview of participants’ real-world coding workflow, consisting of three stages for codebook creation, codebook refinement, codebook application, and different steps taken in inductive and deductive approaches.
[…]
The figure shows a flowchart depicting the 3 stages of the coding workflow: Codebook Creation, Codebook Refinement, and Codebook Application. It has 2 starting points: 1 for inductive, and 2 for deductive, and 1 endpoint labeled Done. 12 elements depict different steps in this process, including visuals representing individual or joint coding. 2 types of lines connect the different elements: arrows depict the element as next step. Dotted arrows depict the next element as optional. The elements representing steps are listed below. Starting point 1: Inductive Step 1: Cluster/Inspect Data. Individual or joint. From this, an arrow leads through pictograms of different data sets to step 2. Step 2: An element with multiple pictograms labeld Individual and 1 labeled Joint are depicted. They are connected to elements labeled Codebook 1 to Codebook N. An arrow leads to the step 3. Step 3. An element labeled First Full Codebook is shown at the end of the stage Codebook Creation. An arrow leads to 3 different elements: Step 3.1, Step 3.2, and Step 3.3. A dotted arrow leads to this arrow, labeled Optional IRR. Step 3.1 (Re-)Apply Codebook in the Stage of Codebook Application. An arrow leads to the next element: Refine Codebook, including labels for Remove Codes and Merge Codes. A dotted arrow labeled IRR high leads to the final step: Done. From this dotted arrow, another arrow with the description IRR low leads back to the element Refine Codebook. Another arrow labeled never ending process leads back to the element (Re-)Apply Codebook. Step 3.2: Test Codebook in the Stage of Codebook Refinement. This element consists of 3 pictograms for coders. 1: Individual, 2: Joint, 3: Joint with new. An arrow leads to step 4. Step 4: Refine Codebook in the Stage of Codebook Refinement. The element contains labels for Remove codes and Reduce granularity. An arrow leads to step 5. Step 5. (Re-)Apply Codebook in the stage of Codebook Application. It contains labels for Rest of data and Seen data. It further shows iterations labeled Single, Multiple, or Random Multiple. A dotted arrow labeled IRR high leads back to the same element (Re-)Apply Codebook. A dotted arrow labeled IRR low leads back to step 4: Refine Codebook. Another dotted arrow leads to the last element: Done. Step 3.3: An element labeled Discuss with Coded Data. From here, an arrow leads to step 5: (Re-)Apply Codebook. Starting point 2: Deductive Step 1: An element labeled Deduce from Literature / Research Questions. A dotted arrow leads to Step 2 in Inductive Coding. A straight arrow leads to the next step in Codebook Application. Step 2.2: An element labeled Apply Codebook. A dotted arrow leads to the last element: Done.
@@
At a high level, our participants’ coding processes consist of three main tasks. (1) The Codebook Creation – here, the very first version of a codebook is developed. (2) The Codebook Refinement – where researchers adjust and update the first codebook version based on discussion, or by applying it to new, unseen data. Lastly, in step (3) , the Codebook Application, the “final” codebook is applied to the rest of the data – however, even with thorough preparation of the codebook, participants reported more codes may be added in this phase. As illustrated in Figure 1, the majority of our participants followed an inductive coding strategy. Only two reported having created their codebooks based on a priori information (deductive coding). Team involvement varied substantially. Some participants coded alone, while most had a lead researcher handle codebook creation and refinement, with team input during discussions. In some cases, new researchers applied the final codebook to the full dataset.
@@
Most participants began by gaining a general overview of the data, skimming through subsets, or clustering them according to certain high-level criteria. Depending on the project, either the entire dataset or a specific subset (often around 10% of the data) was then used for the first version of the codebook. This was done either by (a) one lead researcher, (b) multiple team members who all created their own codebook, which was then combined into one through discussion, or (c) as a joint endeavor where multiple team members created the codebook together. In certain instances, this codebook was then applied to the rest of the data, followed by iterative codebook refinements, referred to as a “never-ending process,” as the codebook application progressed and new themes emerged or began to overlap. However, most reported iteratively applying the codebook to new small sets of data and refining the codebook along the way by removing codes and reducing granularity until all team members were satisfied with the result. Finally, the codebook was applied to the rest of the data. This allowed the determination of a final code count.
@@
There was an almost equal split between those who mentioned that the codebook was only applied by one researcher and those who did multiple codings to be able to calculate inter-rater reliability (IRR). One researcher reported doing random verifications on individual data sets while the full coding had been done by another researcher. IRR scores were calculated at different stages: some did so after creating the first full codebook to check for potential refinements, while others calculated it only at the end for reporting.
@@ 4.1.3. Pains and Gains
Researchers discussed several challenges and difficulties they commonly encounter during QDA processes. These “pain points” provide insights into the complexities of QDA and causes for frustrations. Researchers also highlighted parts of the workflow they enjoyed, which we refer to as “fun points”. Understanding both is crucial to identify opportunities to improve and support QDA workflows and to implement AI support in a way that is desirable for researchers.
@@
Data often needs to be pre-processed, e. g., transcription of audio recordings. Some participants found this a “boring” and “tedious” task, especially when the transcriptions had to be manually corrected or segmented. Several participants expressed particularly enjoying the process of familiarizing themselves with the processed data and finding valuable insights and learning opportunities. This deep engagement was often highlighted as a rewarding aspect of QDA.
@@
    Managing the Codebook. Several participants reported difficulties in managing the codebook’s complexity as it grows, having concerns about losing sight of the “big picture” and ensuring that codes remain coherent. Inferring guiding hierarchies also posed challenges because, it can be “painful to put them into high-level themes without missing an important point.” (P1). This pain point was particularly pronounced among participants who handled large datasets or repetitive coding tasks.
@@
    Restructuring the Codebook Most participants expressed concerns about restructuring the codebook. They dreaded having to go back and adapt codes, update the codebook, or even start anew if the codebook ended up not fitting the data well: ”where the codebook begins is very different from where it ends.“ (P5)
Despite these struggles, several participants acknowledged that the process of developing a codebook is highly educational. Putting codes together and synthesizing themes was also described as exciting and rewarding, offering a sense of accomplishment as patterns and insights begin to emerge.
@@
Participants described codebook application as one of the most time-consuming steps in QDA. They expressed frustration when adjustments to the codebook were necessary later in the process: “When new codes were identified, the codebook had to be updated, and already-coded data had to be revisited to determine if the new code applied.” (P15) Nonetheless, nearly half of the participants perceived the process of applying the codebook to the data as satisfying.
@@
Most participants reported that while collaboration and iterative discussions are essential to their QDA process, those can be tedious and frustrating, particularly in cases where multiple interpretations are possible or different backgrounds are present, making it difficult to reach a consensus. A few participants described onboarding others and initial discussions as cumbersome.
@@
Throughout the QDA process, participants use a variety of tools, some of which they report having difficulty with. For one participant, managing multiple tools was burdensome, as each tool specialized in different tasks, requiring regular data transfers between them. Other complaints about QDA software included challenges with file management, import and export functions, and merging projects when multiple collaborators were involved. Regarding MAXQDA, some participants felt overwhelmed by the wealth of complex features.
@@
Participants with limited QDA experience often expressed concerns about the lack of structure and guidance in QDA. They reported feeling uncertain about whether they were conducting the analysis correctly, expressing a need for assistance, such as having a second person review their work. This uncertainty was exacerbated by participants’ perception of the coding process as messy and subjective, as they mentioned frequently questioning whether their observations had objective value or were merely personal insights.

### v2

@@ 3.1. Study Materials
Our study included a pre-questionnaire and a semi-structured interview with a sketching task. The full interview guide is available in Appendix A. Both the complete pre-questionnaire and the interview guide are available in our anonymized OSF repository.
@@
The interview guide was organized into three parts: (a) participants’ real-life QDA workflows (b) pain points and gain points in QDA (c) perspectives on the potential for AI support
@@
Lastly (c), we explored participants’ expectations around AI-assisted QDA. We inquired about the conditions under which they would be willing to integrate AI into their workflows, which steps they believed could be automated, and which aspects should remain human-driven. Finally, participants were given the opportunity to share any additional thoughts on the topic.
@@ 3.2. Recruitment & Participants
We used purposive and snowball sampling methods (Palinkas et al., 2015) to recruit participants who had conducted at least one study involving QDA through our professional networks within the HCI and Usable Security and Privacy (USP) communities. We approached researchers directly and invited them to refer interested colleagues, leveraging peer identification to build a diverse and qualified sample. Our recruitment strategy was designed to capture variation across research domains and methodological traditions relevant to CSCW and HCI. We categorized participants’ primary research field based on their self-description. To capture a broad range of experiences and perspectives, we took care to recruit participants from diverse group affiliations, different study approaches, topic choices, and educational backgrounds. Our goal was to achieve data saturation regarding participants’ workflows, pain points, and gain points.
@@
Table 1 summarizes the participant sample. We interviewed 15 HCI researchers with backgrounds in computer science, cybersecurity, engineering, and social sciences, all firmly embedded in HCI research, with some specifying to focus on USP. Twelve participants were early career researchers (mostly PhD candidates), and three were senior researchers. Eight had worked on 1-3 QDA projects, four on 4-6, and three on more than 7 projects, with qualitative experience ranging from 1 to 9 years. We recruited participants until data saturation was reached. We reached data saturation around interview 12 and conducted three additional interviews to confirm this.
Table 1. Detailed information on our interview participants’ research backgrounds, their QDA experience, and their primary QDA tools and methods. Field categorizations (HCI, USP, Security) are based on participants’ self-description.
@@
Participants first completed the pre-questionnaire, which included an overview of the study procedure and a consent form. Afterward, we scheduled interviews. The interviews were semi-structured, conducted face-to-face, and lasted approximately 45 minutes. Interviews were conducted in either English ($n=9$) or German ($n=6$), depending on participants’ native language and preference. Two researchers were present: a lead interviewer who conducted the session and took notes, and a second researcher who created detailed protocols, including occasional verbatim quotes. To protect participant privacy and follow data minimization principles, we chose not to record audio. The lead interviewer, fluent in both German and English, conducted all interviews for consistency. The second researcher varied among four researchers and was either fluent in English or fluent in both languages as needed. To ensure objectivity, they noted down points that stood out without performing any immediate analysis or interpretation. The researchers then compared, merged, and verified the completeness of their notes, ensuring no conflicts arose. Finally, when necessary, the researchers manually translated their synthesized interview protocols into English for subsequent data analysis.
We piloted our study with three participants to evaluate the procedure and interview flow. As no changes were made, these pilot sessions were included in the final dataset, resulting in 15 participants.
@@ 3.3. Ethical Considerations & Data Protection
The study was conducted in accordance with the Menlo Report’s ethical principles (Kenneally and Dittrich, 2012) and approved by our institution’s ethics committee. All participants gave informed consent via the pre-questionnaire and again at the start of the interview. Before the interviews began, the interviewer walked the participants through the consent form again and gave them the opportunity to ask questions and asked for their consent again. To protect the privacy of the participants, they were given pseudonyms after completing the questionnaire and all data was pseudonymized. We stored and managed all data according to national data protection regulations and in accordance with our institution’s policies on confidentiality and data protection.
@@ 3.4. Data Analysis
Three authors developed an initial set of deductive codes for tools and resources, data types, and participants’ overall willingness to use AI, based on the pre-questionnaire answer options and interview guide. They each coded three protocols for the coding process, pain points, fun points, and concerns and conditions for using AI within QDA to derive common themes. This process involved closely examining the protocols to identify and conceptualize subcategories. All three coders then created the final codebooks by discussing the identified themes and resolving conflicts in a collaborative session. Two authors then coded the entire dataset, resolving disagreements through discussion until full agreement was reached. To help us later illustrate our findings, the coders included example quotations from the protocols in our codebooks, both paraphrased and verbatim.
Additionally, we analyzed the drawings from the sketching task to derive commonalities and differences in the coding process and extracted participants’ suggestions for concrete AI applications as in vivo codes. Two coders reviewed sketches together with protocols to create a combined metaview of current QDA practices. Each participant’s workflow was redrawn and sequentially mapped into a shared metaview. New elements were added as the emerged, recurring ones were skipped. Closely related processes were later merged to reduce complexity. We differentiated elements taking into account the SeeMe-method developed by Herrmann (Herrmann, 2012). SeeMe is a semi-structured socio-technical modeling technique used for task analysis that involves three visualization elements: (1) roles assigned to individuals or teams, (2) activities that are of a more dynamic characteristic and represent change functions, and (3) entities, defined as objects from the physical world. For entities, we excluded tools and software support from the metaview as they were heavily dependent on the participant’s individual preference, but generally involved in each process element.
@@ 4. Results
Here, we present HCI researchers’ QDA workflows, including their pain points (Section 4.1) and report participants’ willingness to use AI support along with concerns and conditions (Section 4.2). In Appendix B, we provide the complete codebooks from our interviews, including descriptions, examples, and quantities for each code. Appendix C shows example workflow sketches.
@@ 4.1.1. Contexts of Studied QDA Workflows
In the pre-questionnaire, participants reported having analyzed open-ended survey responses ($n=12$), interview transcripts ($n=11$), written documents ($n=8$), and social media posts ($n=5$) as the most common data types. Most participants ($n=9$) reported using more than one tool for qualitative coding, including MAXQDA ($n=12$), Excel ($n=6$), and Google Docs ($n=4$). Participants also indicated experience with multiple methodological frameworks, often adapting them based on the nature of the data and their research context. Thematic Analysis ($n=11$), Grounded Theory ($n=6$), and Content Analysis ($n=4$) were mentioned most frequently.
@@ 4.1.2. Uncovering Participants’ Actual Coding Workflow
Figure 1. Detailed overview of participants’ real-world coding workflow, consisting of three stages for codebook creation, codebook refinement, codebook application, and different steps taken in inductive and deductive approaches.:
@@
At a high level, our participants’ coding processes generally followed three main steps: (1) Codebook Creation– developing an initial version of the codebook, (2) Codebook Refinement– updating the codebook through discussion, or by applying it iteratively to new, unseen data, and (3) Codebook Application– applying the codebook to the full dataset, with occasional ongoing adjustments. As illustrated in Figure 1, the majority of our participants followed an inductive approach. Only two created their codebooks based on a priori information with deductive coding. Team structures varied: some participants coded alone, while most had a lead researcher handle codebook creation and refinement, with team input during discussions. In some cases, new researchers applied the final codebook to the full dataset.
@@
Typically, participants began by reading through a subset of the data to develop an overview. Depending on the project, either the entire dataset or a specific subset (often around 10% of the data) was then used to develop an initial codebook. This was done either individually, in parallel by multiple coders who later merged their versions, or collaboratively from the start. In certain instances, this codebook was then applied to the rest of the data, followed by iterative codebook refinements, referred to as a “never-ending process,” as the codebook application progressed and new themes emerged or began to overlap. However, most reported iteratively applying the codebook to new small sets of data and refining the codebook along the way by removing codes and reducing granularity until all team members were satisfied with the result. Finally, the codebook was applied to the rest of the data. This allowed the determination of a final code count.
@@
There was an almost equal split between those who mentioned that the codebook was only applied by one researcher and those who did multiple codings to be able to calculate inter-rater reliability (IRR). In some cases, IRR was used during early codebook development to support refinement; in others, it was calculated only at the end for reporting purposes. One participant described using random checks for verification, rather than full double-coding.
@@ 4.1.3. Pains and Gains
Researchers discussed several challenges and difficulties (“pain points”) and rewarding aspects (“fun points”) of their QDA workflows. Understanding both helps identify opportunities for AI support that align with researchers’ values and practices.
@@
Data often needs to be pre-processed, e. g., transcription of audio recordings. Some participants found this a boring and tedious task, especially when the transcriptions had to be manually corrected or segmented. Several participants expressed particularly enjoying the process of familiarizing themselves with the processed data and finding valuable insights and learning opportunities. This deep engagement was often highlighted as a rewarding aspect of QDA.
@@
    Managing the Codebook. Several participants reported difficulties in managing the codebook’s complexity as it grows, having concerns about losing sight of the big picture and ensuring that codes remain coherent. Inferring hierarchies also posed challenges because it can be “painful to put them into high-level themes without missing an important point.” (P1). This pain point was particularly pronounced among participants who handled large datasets or repetitive coding tasks.
@@
    Restructuring the Codebook. Most participants expressed concerns about restructuring the codebook. They dreaded having to go back and adapt codes, update the codebook, or even start anew if the codebook ended up not fitting the data well: ”where the codebook begins is very different from where it ends.“ (P5)
Despite these struggles, several participants acknowledged that the process of developing a codebook is highly educational. Synthesizing patterns and refining themes was also described as exciting and rewarding, offering a sense of accomplishment.
@@
Participants described codebook application as one of the most time-consuming tasks. They expressed frustration when adjustments to the codebook were necessary later in the process: “When new codes were identified, the codebook had to be updated, and already-coded data had to be revisited to determine if the new code applied.” (P15) Nonetheless, nearly half of the participants perceived the process of applying the codebook to the data as satisfying.
@@
Most participants reported that while collaboration and iterative discussions are essential to their QDA process, those can be tedious and frustrating, particularly in cases where multiple interpretations are possible or different backgrounds are present, making it difficult to reach consensus. A few participants described onboarding others and initial discussions as cumbersome.
@@
Throughout the QDA process, participants use a variety of tools, some of which they report having difficulty with. For one participant, managing multiple tools was burdensome, as each tool specialized in different tasks, requiring regular data transfers between them. Issues included cumbersome data transfers, file management, import/export limitations, and collaboration features. In particular, MAXQDA’s complexity overwhelmed some users.
@@
Participants with limited QDA experience often expressed concerns about the lack of structure and guidance in QDA. They reported feeling uncertain about whether they were conducting the analysis correctly, expressing a need for assistance, such as having a second person review their work. This was linked to the perception that coding is subjective and inherently messy, making it difficult to evaluate whether their observations had objective value or were merely personal insights.